# LPR: Large Language Models-Aided Program Reduction

### v2

@@ 3.3. Transformations
To further search for more reduction opportunities on a bug-triggering program via the LLM, we propose five general transformations to guide the LLM, i.e., Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification and Variable Elimination.
Function Inlining. This transformation identifies a function and performs function inlining to eliminate all call sites of this function, and instead substitutes them with corresponding the function body. As functions are prevalent presented in bug-triggering programs, there is significant room for function inlining to reduce tokens or provide further reduction opportunities.
Loop Unrolling. Loop unrolling, also known as loop unwinding, is a widely used loop transformation approach to optimize the execution. In this task, it can also be employed to find more reduction opportunities. Loop Unrolling identifies a loop structure and attempts to unroll the loop into a code snippet repeating a single iteration. Motivation lies in that language-generic reducers may be incapable of dismantling or directly removing the loop structure, while they may be able to reduce the repeated code after loop unrolling, as shown in Figure 1.
@@ 4. Evaluation
In this section, we evaluate the effectiveness and efficiency of LPR. Specifically, we conducted the following research questions.
RQ1. What is the effectiveness of LPR in program reduction?
RQ2. What is the efficiency of LPR in program reduction?
RQ3. What is the effectiveness of each transformation in LPR?
@@ 4.1. Experimental Setup
Within the workflow of LPR, we employ Perses (Sun et al., 2018) as the language-agnostic reducer due to its superior efficiency compared to Vulcan. Additionally, we utilize OpenAI API (ope, 2023a), specifically the gpt-3.5-turbo-0613 version, to serve as the LLM. We also develop a variant named LPR+Vulcan, which invokes Vulcan to further reduce the program after LPR finishes. For a fair comparison, all algorithms were executed in a single-process, single-thread environment.
Benchmarks. To measure the effectiveness and efficiency of LPR across various languages, we employ three benchmark suites: Benchmark-C, Benchmark-Rust and Benchmark-JS. The Benchmark-C, previously collected and utilized by previous studies (Xu et al., 2023; Tian et al., 2023c; Zhang et al., 2023), comprises 20 large complex programs triggering real-world bugs in LLVM or GCC. Benchmark-Rust, incorporating 20 bug-triggering Rust programs, has also been used in prior research (Xu et al., 2023). We further craft Benchmark-JS, a non-public benchmark suite, for this study. Specifically, we use FuzzJIT (Wang et al., 2023) to fuzz a prevalent JavaScript engine, i.e., JavaScriptCore (version c6a5bcc), and then randomly collect 10 programs that cause miscompilations in JIT compiler. Since the programs and reduced programs in Benchmark-JS are not publicly available and thus not in the training sets of LLMs. The evaluation on Benchmark-JS helps us investigate whether LPR suffers from the data leakage problem (Wu et al., 2023). In total, the evaluation benchmarks encompass 50 programs triggering real-world bugs in compilers, spanning across three popular programming languages.
@@
Configuration. If not otherwise specified, our experiments are conducted by invoking OpenAI API (version gpt-3.5-turbo-0613), with the proposed multi-level prompt and transformations in section 3.3 and section 3.4. To effectively harness the inherent randomness of LLMs, we set temperature=1.0. This high value encourages the LLM to generate more diverse outcomes (ope, 2023c). Additionally, we employ n=5 to generate five distinct results for every query (ope, 2023b), enabling us to choose the smallest passing program as the optimal result. All the rest configurations are set to their default values.
4.2. RQ1: Effectiveness
We measure the effectiveness of LPR, LPR+Vulcan and baseline algorithms via the final program size in token. A smaller size is favored, as it signifies the removal of more bug-irrelevant code, thereby saving developers more manual effort. The effectiveness of each algorithm on all three benchmark suites is presented in Table 1. Due to the randomness of LPR, we repeat five times for LPR and LPR+Vulcan on every benchmark, and display the mean and standard deviation in the table. On each benchmark, the minimal results are highlighted in bold.
@@ 4.2. RQ1: Effectiveness
Further analysis into these benchmarks reveals that Vulcan and transformations in LPR are complementary on Benchmark-Rust. Vulcan shows higher effectiveness on reducing relatively smaller programs, as evidenced by the average original size of 43 tokens in the 9 programs where it is better than LPR. In contrast, the 15 programs where LPR is proved better than Vulcan have an average of 269 tokens, indicating its proficiency on reducing relatively larger programs. Our speculation is that Vulcan and LPR target different reduction opportunities. Vulcan performs identifier/sub-tree replacement and local exhaustive search. Such reducers, while lacking in semantic analysis, find reduction opportunities in a ”brute-force” manner and is particularly effective at uncovering less obvious reduction opportunities. On the other hand, LPR employs more semantic and intelligent transformations, adeptly and systematically analyzing and reducing a complex program step by step. Moreover, results of LPR+Vulcan in the last column prove the complementary characteristic between LPR and Vulcan, which achieve the best in 16 out of the total 20 benchmarks.
@@
4.3. RQ2: Efficiency
In this research question, we measure the time consumed by each technique on three benchmarks. Shorter time indicates higher efficiency. Table 2 shows the results. Since both Vulcan and LPR perform reduction on top of Perses’s results, it is impossible for these two algorithms to take less time than Perses. Besides, as a highly efficient tree-based reduction algorithm, Perses is generally faster than C-Reduce. Therefore, we focus on the comparison among Vulcan, C-Reduce and LPR.
@@ 4.3. RQ2: Efficiency
In the Benchmark-C, compared to Vulcan and LPR, C-Reduce generally has a shorter reduction time. This is expected, as C-Reduce’s transformations are specifically designed for C languages, whereas Vulcan approaches the problem in a more unguided and brute-force manner, and LPR’s attempts are not specifically designed for C and rely more on the efficiency of LLMs. On average, LPR takes 1:54:54, which is % shorter than 2:08:46 of Vulcan. However, in terms of the average percentage difference in time, LPR requires 45.83% more time compared to Vulcan in Benchmark-C. The main reason of such a result is that LPR is more efficient than Vulcan when the program is large and complex, as the transformations it performs are aware of the semantics and have a higher success rate in reducing the program. However, when the program is small and simple, Vulcan can finish quickly since its search space becomes considerably small, but for LPR, the time consumed by the LLM is not decreased significantly and becomes dominant, which makes LPR less efficient than Vulcan in these benchmarks. On Benchmark-Rust, The results indicate a similar trend, i.e., LPR tends to be more efficient when reducing complex programs while Vulcan is more efficient in small and simple benchmarks. If we only keep benchmarks where both tools take longer than one hour, LPR requires 4.15% and 21.69% less time compared to Vulcan.
In our analysis of Benchmark-Rust, we observed that both LPR and Vulcan can consume an extreme long time on certain benchmarks. For instance, Vulcan requires 20 hours for Rust-99830, while LPR takes a similar duration on Rust-66851 in a specific run. This extensive time consumption is often due to the frequent invocation of Perses, which only achieves marginal progress with each transformation. The prolonged duration of Perses is primarily attributed to the strict syntax of the Rust language. Considering that program reduction is an NP-complete problem, this inefficiency might be optimized in future, whereas it cannot be completely eliminated.
After further in-depth analysis, another interesting fact emerges. On the three benchmark suites, the average time taken by LPR is 1.915, 1.839, and 0.174 hours, respectively. However, within these durations, the time spent waiting for the LLM responses accounts for %, %, and %. This experiment involves invoking the OpenAI API, and might have been limited by high user traffic and few computational resources allocated. Considering the ongoing advancements in LLMs technology, we believe that LPR’s efficiency will be substantially improved in the future.
@@ 4.4. RQ3: Effectiveness of Each Transformation
To answer this question, we delve into the impact of each transformation on program size, alongside their potential to escape the local minimal program and unlock new reduction opportunities. Our in-depth analysis focuses on Benchmark-C because of the complexity of bug-triggering programs.
@@
For each transformation, we measure how the program size change in each benchmark in Benchmark-C after a specific transformation is performed, and plot the size changes into a box-plot and red dots, as shown in Figure 4. Additionally, since each transformation is immediately followed by an invocation of Perses, we also monitor the cumulative size changes resulting from both the transformation and its subsequent Perses reduction. These changes are depicted by the blue dots in the second box-plot of each subplot in Figure 4. Given that a transformation is performed in every iteration, we compute the average size change by dividing the total sum of size changes for that transformation by the number of iterations. This approach enables us to thoroughly comprehend how each transformation influences program size and assess its capacity to provide further reduction opportunities to Perses.
@@
According to size changes induced by each transformation alone, i.e., the left box-plot with red dots, we can find two trends. First, Function Inlining, Data Type Elimination and Variable Elimination are more likely to reduce the program size by themselves, with an average of size change -13.2, -4.7 and -4.3, separately. However, for the rest two transformations, i.e., Loop Unrolling and Data Type Simplification, most of the program sizes increase instead, with an average of 14.9 and 1.3 respectively. This is expected, as such transformations will generally transform the program into a larger one. Loop Unrolling, disassembles a loop into repeated lines of code, and leads to size increase temporarily. Data Type Simplification can dismantle a variable in complex data type, e.g., structure, into a list of members in primary data types, which may requires more tokens to declare and initialize each variable.
@@
For size changes induced by both a transformation and the subsequent execution of Perses, i.e., the right box-plot with blue dots, all of the proposed transformation result in size decreases. The fact that the right box-plot is generally lower than the left one indicates that Perses often further removes tokens after the transformation is applied. For Loop Unrolling and Data Type Simplification, even though they usually introduce more tokens to the program, they expose reduction opportunities for the following execution of Perses, and eventually result in a smaller result.
@@
To further understand the impact of each transformation on the entirety of Benchmark-C, we provide a detailed analysis of their contributions. Specifically, on the 5 repeated experiments on 20 benchmarks, we calculate the average size reduction brought about by each transformation across all these 100 runs by summing up the size decreases attributed to the transformation in each benchmark and then computing the mean. This is illustrated in Figure 5. Additionally, we quantify the prevalence of each transformation by counting the number of benchmarks in which it induces a size decrease, as demonstrated in Figure 6. These metrics together offer a comprehensive view of how each transformation influences file sizes and how frequently they take effect within Benchmark-C.
@@
From Figure 5, it is evident that every transformation contributes to further size reduction in Benchmark-C. Specifically, while Function Inlining is responsible for a reduction of 88.2 tokens on average, contributing 62.12% to the overall decrease, Loop Unrolling shows the minimal impact, contributing only 4.8 tokens, 3.35% to the overall decrease. This highlights the varying degrees of influence each transformation has on code size. Further insights from Figure 6 reveal that Data Type Elimination affects all benchmarks, likely due to the ubiquitous presence of typedef across all benchmarks. In contrast, Loop Unrolling is the least prevalent transformation, affecting merely 6 on average out of 20 benchmarks. This outcome is expected, considering that not all programs involve loop structures, and not every loop is irrelevant to the bugs in compilers. Besides, the relatively higher standard deviations observed in the Data Type Simplification and Loop Unrolling suggest that these transformations present greater challenges for the LLM to manage effectively.
@@ 5. Discussion
In this section, we discuss the effectiveness of multi-level prompts and, the performance of the LLM under different temperatures.
@@ 5.2. The Impact of Temperature
In our experiments, we consistently set the temperature parameter of LLMs to 1.0. To measure how this parameter affects the performance, we rerun experiments under multiple temperatures, i.e., 0.75, 0.5, 0.25, 0. Note that higher temperature instructs the LLM to generate more creative and diverse results. Due to limited resources and time, we evaluate on 10 benchmarks in Benchmark-C that demonstrated the fastest completion times under the default configuration.
Perses | Vulcan | LPR with gpt-3.5-turbo-0613 | |||||
Mean | 161.4 | 102.8 |

### v3

@@ 3.3. Transformations
To further search for reduction opportunities on a bug-triggering program via the LLM, we propose five general transformations to guide the LLM, i.e., Function Inlining, Loop Unrolling, Data Type Elimination, Data Type Simplification and Variable Elimination.
Function Inlining. This transformation identifies a function and inlines it to eliminate all call sites of this function, and instead substitutes them with the corresponding function body. As functions are commonly found in bug-triggering programs, there is significant room for function inlining to reduce tokens or provide further reduction opportunities.
Loop Unrolling. Loop unrolling, also known as loop unwinding, is a widely used loop transformation approach to optimize the execution. In this task, it can also be employed to find more reduction opportunities. Loop Unrolling identifies a loop structure and attempts to unroll the loop into a code snippet repeating a single iteration. This is motivated by the fact that language-generic reducers may be incapable of dismantling or directly removing the loop structure, while they may be able to reduce the repeated code after loop unrolling, as shown in Figure 1.
@@ 4. Evaluation
In this section, we evaluate the reduction effectiveness and efficiency of LPR. Specifically, we conducted the following research questions.
- RQ1:
What is the effectiveness of LPR in program reduction?
- RQ2:
To what extent is the efficiency of LPR in program reduction perceived by users?
- RQ3:
What is the effectiveness of each transformation in LPR?
@@ 4.1. Experimental Setup
Within the workflow of LPR, we employ Perses (Sun et al., 2018) as the language-agnostic reducer due to its superior efficiency compared to Vulcan. Additionally, we utilize OpenAI API (ope, 2023a), specifically the gpt-3.5-turbo-0613 version, to serve as the LLM. We also develop a variant named LPR+Vulcan, which invokes Vulcan to further reduce the program after LPR finishes. The experiments are conducted on an Ubuntu 22 server with an Intel(R) Xeon(R) Gold 6348 CPU @ 2.60GHz and 512 GB RAM. For a fair comparison, all algorithms were executed in a single-process, single-thread environment.
Benchmarks. To measure the effectiveness and efficiency of LPR across various languages, we employ three benchmark suites: Benchmark-C, Benchmark-Rust and Benchmark-JS. The Benchmark-C, previously collected and utilized by previous studies (Xu et al., 2023; Tian et al., 2023c; Zhang et al., 2023), comprises 20 large complex programs triggering real-world bugs in LLVM or GCC. Benchmark-Rust, incorporating 20 bug-triggering Rust programs, has also been used in prior research (Xu et al., 2023). We further craft Benchmark-JS, a non-public benchmark suite, for this study. Specifically, we use FuzzJIT (Wang et al., 2023b) to fuzz a prevalent JavaScript engine, i.e., JavaScriptCore (version c6a5bcc), and then randomly collect 10 programs that cause miscompilations in JIT compiler. Since the programs and reduced programs in Benchmark-JS are not publicly accessible at the time of LLMs’ training, and thus not in the training sets of LLMs. The evaluation on Benchmark-JS helps to investigate whether LPR suffers from the data leakage problem (Wu et al., 2023). In total, the evaluation benchmarks encompass 50 programs triggering real-world bugs in compilers, spanning across three popular programming languages.
@@
Configuration. If not otherwise specified, our experiments are conducted by invoking OpenAI API (version gpt-3.5-turbo-0613), with the proposed multi-level prompt and transformations in section 3.3 and section 3.4. To effectively harness the inherent randomness of LLMs, we set temperature=1.0. This high value encourages the LLM to generate more diverse outcomes (ope, 2023c). Additionally, we employ n=5 to generate five distinct results for every query (ope, 2023b), enabling us to choose the smallest program preserving the property as the optimal result. All the rest configurations are set to their default values.
4.2. RQ1: Reduction Effectiveness
We measure the effectiveness of LPR, LPR+Vulcan and baseline algorithms via the final program sizes in tokens. A smaller size is favored, as it signifies the removal of more bug-irrelevant code, thereby saving developers more manual effort. The effectiveness of each algorithm on all three benchmark suites is presented in Table 1. Due to the randomness of LPR, we repeat five times for LPR and LPR+Vulcan on every benchmark, and display the mean value and standard deviation value in the table. On each benchmark, the minimal results are highlighted in bold.
@@ 4.2. RQ1: Reduction Effectiveness
Further analysis into these benchmarks reveals that Vulcan and transformations in LPR are complementary on Benchmark-Rust. Vulcan demonstrates greater effectiveness in reducing programs of relatively smaller size, as highlighted by the average original size of 43 tokens across 9 programs, outperforming LPR. In contrast, the 15 programs where LPR is proved better than Vulcan have an average of 269 tokens, indicating its proficiency in reducing relatively larger programs. Our speculation is that Vulcan and LPR target different reduction opportunities. Vulcan performs identifier/sub-tree replacement and local exhaustive search. Such reducers, while lacking in semantic analysis, find reduction opportunities in a ”brute-force” manner and are particularly effective at uncovering less obvious reduction opportunities. On the other hand, LPR employs more semantic and complex transformations, adeptly and systematically analyzing and reducing a complicated program step by step. Moreover, the results of LPR+Vulcan in the last column prove the complementary characteristic between LPR and Vulcan, which achieve the best in 16 out of the total 20 benchmarks.
@@
4.3. RQ2: Reduction Efficiency
In this research question, we measure the time elapsed when users employ each technique for program reduction.111 Please note that the time measured in this research question cannot be directly used to compare the computational resources consumed by LPR with those consumed by other techniques. LPR is standing on the shoulders of giants, i.e., LLMs, of which computations heavily rely on GPUs, and it is infeasible for us to measure the actual resource consumption of each query to OpenAI API. In contrast, Perses, Vulcan, and C-Reduce only leverage CPUs. Shorter time indicates higher efficiency, and Table 2 shows the results. Since both Vulcan and LPR perform reduction on top of Perses’s results, it is impossible for these two algorithms to take less time than Perses. In addition, as a highly efficient tree-based reduction algorithm, Perses is generally faster than C-Reduce. Therefore, we focus on the comparison among Vulcan, C-Reduce and LPR.
@@ 4.3. RQ2: Reduction Efficiency
In Benchmark-C, compared to Vulcan and LPR, C-Reduce generally has a shorter reduction time. This is expected, as C-Reduce’s transformations are specifically designed for C languages, whereas Vulcan approaches the problem in a more unguided and brute-force manner, and LPR’s attempts are not specifically designed for C and rely more on the efficiency of LLMs. On average, LPR takes 1h:54m:54s, which is % shorter than 2h:08m:46s of Vulcan. However, in terms of the average percentage difference of each benchmark, LPR requires 45.83% more time compared to Vulcan in Benchmark-C. The main reason for such a result is that LPR is more efficient than Vulcan when the program is large and complex, as the transformations it performs are aware of the semantics and have a higher success rate in reducing the program. However, when the program is small and simple, Vulcan can finish quickly since its search space becomes considerably small, but for LPR, the time consumed by the LLM is not decreased significantly and becomes dominant, making LPR less efficient than Vulcan in these benchmarks. On Benchmark-Rust, The results indicate a similar trend, i.e., LPR tends to be more efficient when reducing complex programs while Vulcan is more efficient in small and simple benchmarks. If we only keep benchmarks where both tools take longer than one hour, LPR requires 4.15% and 21.69% less time compared to Vulcan.
In our analysis of Benchmark-Rust, we observed that both LPR and Vulcan can consume an extremely long time on certain benchmarks. For instance, Vulcan requires 20 hours for Rust-99830, while LPR takes a similar duration on Rust-66851 in a specific run. This extensive time consumption is often due to the frequent invocation of Perses, which only achieves marginal progress with each transformation. The prolonged duration of Perses is primarily attributed to the strict syntax of the Rust language. Considering that program reduction is an NP-complete problem, this inefficiency might be optimized in the future, whereas it cannot be completely eliminated.
After further in-depth analysis, another interesting fact emerges. On the three benchmark suites, the average time taken by LPR is 1.915, 1.839, and 0.174 hours, respectively. However, within these durations, the time spent waiting for the LLM responses accounts for %, %, and %. Additionally, we measure the efficiency from more perspectives. Take Benchmark-C as an example: the average expense to finish each benchmark amounts to $0.42, with each benchmark requiring 41 queries on average, and every query consuming approximately 39 seconds on average. This experiment involves invoking the OpenAI API, and might have been limited by high user traffic and few computational resources allocated. Considering the ongoing advancements in LLMs technology, we believe that LPR’s efficiency will be substantially improved in the future.
@@ 4.4. RQ3: Effectiveness of Each Transformation
To answer this question, we delve into the impact of each transformation on program size, alongside their potential to escape the local minimal program and unlock new reduction opportunities. Our in-depth analysis focuses on Benchmark-C because of its complexity of bug-triggering programs compared to other benchmark suites, ensuring enough room for each transformation to take effect.
@@
For each transformation, we measure how the program size changes in each benchmark of Benchmark-C after a specific transformation is performed, and plot the size changes into a box-plot and red dots, as shown in Figure 4. Additionally, since each transformation is immediately followed by an invocation of Perses, we also monitor the cumulative size changes resulting from both the transformation and its subsequent Perses reduction. These changes are depicted by the blue dots in the second box-plot of each subplot in Figure 4. Given that a transformation is performed in every iteration, we compute the average size change by dividing the total sum of size changes for that transformation by the number of iterations. This approach enables us to thoroughly comprehend how each transformation influences program size and assess its capacity to provide further reduction opportunities to Perses.
@@
According to size changes induced by each transformation alone, i.e., the left box-plot with red dots, we can find two trends. First, Function Inlining, Data Type Elimination and Variable Elimination are more likely to reduce the program size by themselves, with an average size change -13.2, -4.7 and -4.3, separately. However, for the rest two transformations, i.e., Loop Unrolling and Data Type Simplification, most of the program sizes increase instead, with an average of 14.9 and 1.3 respectively. This is expected, as such transformations will generally transform the program into a larger one. Loop Unrolling, disassembles a loop into repeated lines of code, and leads to size increase temporarily. Data Type Simplification can dismantle a variable in a complex data type, e.g., structure, into a list of members in primary data types, which may require more tokens to declare and initialize each variable.
@@
For size changes induced by both a transformation and the subsequent execution of Perses, i.e., the right box-plot with blue dots, all of the proposed transformation result in size decreases. The fact that the right box-plot is generally lower than the left one indicates that Perses often further removes tokens after the transformation is applied. For Loop Unrolling and Data Type Simplification, even though they usually introduce more tokens to the program, they expose reduction opportunities for the following execution of Perses, and eventually result in a smaller program.
@@
To further understand the impact of each transformation on the entirety of Benchmark-C, we provide a detailed analysis of their contributions. Specifically, on the 5 repeated experiments on 20 benchmarks, we calculate the average size reduction brought about by each transformation across all these 100 runs by summing up the size decreases attributed to the transformation in each benchmark and then computing the mean value, as illustrated in Figure 5. Additionally, we quantify the prevalence of each transformation by counting the number of benchmarks in which it induces a size decrease, as demonstrated in Figure 6. The above evaluations together offer a comprehensive view of how each transformation influences program sizes and how frequently they take effect within Benchmark-C.
@@
From Figure 5, it is evident that every transformation contributes to further size reduction in Benchmark-C. Specifically, while Function Inlining is responsible for a reduction of 88.2 tokens on average, contributing 62.12% to the overall decrease. Loop Unrolling shows a minimal effect, contributing only 4.8 tokens, 3.35% to the overall decrease. This highlights the varying degrees of influence each transformation has on program size. Further insights from Figure 6 reveal that Data Type Elimination affects all benchmarks, likely due to the ubiquitous presence of typedef across all benchmarks. In contrast, Loop Unrolling is the least prevalent transformation, affecting merely 6 on average out of 20 benchmarks. This outcome is expected, considering that not all programs involve loop structures, and not every loop is irrelevant to the compiler bugs. In addition, the relatively higher standard deviations observed in the Data Type Simplification and Loop Unrolling suggest that these transformations present greater challenges for the LLM to effectively handle.
@@ 5. Discussion
In this section, we discuss the effectiveness of multi-level prompts the performance of the LLM under different temperatures, and the failures in LPR.
@@ 5.2. The Impact of Temperature
In our experiments, we consistently set the temperature parameter of LLMs to 1.0. To measure how this parameter affects the performance, we rerun experiments under multiple temperatures, i.e., 0.75, 0.5, 0.25, 0. Note that higher temperature instructs the LLM to generate more creative and diverse results. Due to limited resources and time, we evaluate on 10 benchmarks in Benchmark-C with the fastest completion times under the default configuration.
| Perses | Vulcan | LPR | |||||
| Mean | 161.4 | 102.8 | |||||
@@
5.3. Failures in LPR
In LPR, the generation of variants failing to pass the property test is common and acceptable. There are two scenarios where LPR fails to produce variants passing the property test. First, the LLM is generally non-deterministic, it may fail to perform the correct transformation when the program is complex. Therefore, multiple responses are requested in each LLM query to mitigate the impact of potential failures. In another scenario, the transformation may eliminate the bug-triggering pattern, e.g., the compiler crashes on a function call, but the transformation inlines this function. Even if such a transformation is semantically correct, the property test will still fail. When the transformed program is bound to fail, LPR will proceed with the next step using the original program.
Failures indeed lead to more property tests before making any progress, but they are inevitable, as program reduction is a trial-and-error process. Even without LLM, property checks issued by Perses and Vulcan usually have a failure rate of around 90%. As shown in Table 1 and Table 2, LPR is effective and efficient enough, illustrating that failures can be mitigated via multiple responses.